### v1

# A unified approach for degree bound estimates of linear differential operators

Abstract.
We identify a common scheme in several existing algorithms adressing computational problems on linear differential equations with polynomial coefficients. These algorithms reduce to computing a linear relation between vectors obtained as iterates of a simple differential operator known as pseudo-linear map.
We focus on establishing precise degree bounds on the output of this class of algorithms. It turns out that in all known instances (least common left multiple, symmetric product,…), the bounds that are derived from the linear algebra step using Cramer’s rule are pessimistic. The gap with the behaviour observed in practice is often of one order of magnitude, and better bounds are sometimes known and derived from ad hoc methods and independent arguments. We propose a unified approach for proving output degree bounds for all instances of the class at once. The main technical tools come from the theory of realisations of matrices of rational functions and their determinantal denominators.
$D$-finiteness, Pseudo-linear map, Degree bound
journalyear: 2025conference: International Symposium on Symbolic and Algebraic Computation; 2025; Guanajuato, Mexicobooktitle: International Symposium on Symbolic and Algebraic Computation (ISSAC ’25), 2025, Guanajuato, Mexicoccs: Computing Methodologies Algebraic algorithmsccs: Theory of computation Analysis of algorithms
1. Introduction
In this work, we study four central algorithmic problems involving D-finite functions (stanley1980dfinite, ; Kauers23, ), i.e. defined by linear differential equations with polynomial coefficients. We specifically focus on the Hermite reduction-based creative telescoping algorithm for the simple integration of bivariate rational functions (bostan2010complexity, ), the computation of a differential equation for an algebraic function (bostan2007algeqtodiffeq, ) and two closure operations on $D$-finite functions: least common left multiple (vanderHoeven2016complexity_skew_arith, ; bostan2012LCLM, ) and symmetric product (kauers2014bounds_Dfinite_closure, ; stanley1980dfinite, ). These four problems can all be seen as particular instances of a large class of algorithmic problems that share the common property to be expressible with the following algebraic formulation. Let $k$ be a field of characteristic 0 and $\partial_{x}$ be the usual derivative on the rational function field $k(x)$. We are given a sequence $(a_{i})_{i\geq 0}$ of $k(x)^{n}$ defined by the simple inductive formula:
(1) $a_{0}\in k[x]^{n},\quad a_{i+1}=\left(\partial_{x}+T\right)\cdot a_{i},~{}i\geq 0,$
where $T\in k(x)^{n\times n}$ is a $k(x)$-linear map. The problem is to compute the coefficients $\eta_{i}\in k[x]$ of a minimal linear relation $\eta_{\rho}(x)\cdot a_{\rho}+\cdots+\eta_{0}(x)\cdot a_{0}=0$. In (1), the operator $\theta=\partial_{x}+T$ is an example of a pseudo-linear map (jacobson1937pseudolineartransformations, ; bronstein1996introduction, ), and we have $a_{i}=\theta^{i}a_{0}$. In other words, we consider the following problem.
Problem 1.
Given $T\in k(x)^{n\times n}$, $a\in k[x]^{n}$ and letting $\theta=\partial_{x}+T$ and $\rho=\dim\operatorname{span}_{k(x)}(\theta^{i}a,i\geq 0)\leq n$, find $\eta=(\eta_{0},\dots,\eta_{\rho})\in k[x]^{\rho+1}\setminus\{0\}$ such that
(2) $\eta_{\rho}\cdot\theta^{\rho}a+\cdots+\eta_{1}\cdot\theta a+\eta_{0}\cdot a=0.$
In the case $\rho=n$, i.e. $a$ is a cyclic vector for $T$ (churchill2002cyclicvectors, ; bostan2013cyclic, ; adjamagbo1988cyclic, ; barkatou1993algo_uncoupling, ), the above relation gives a characteristic polynomial of $\theta$ (amitsur1954diffpoly_divalgebra, ). Problem 1 can be seen as an extension of the method of Krylov iterates for computing the characteristic polynomial (neiger_pernet_villard_2024_krylov_iterates, ; keller-gehrig_1985_charpoly, ) in the differential context. As further developed in Section 2, each of the four problems mentioned above reduces to solving the linear system (2) with a specific choice of $T$ and $a$.
Hermite reduction-based creative telescoping
[…]
Differential equation for an algebraic function
A power series $\alpha(x)$ over the field $k$ is said to be algebraic if it is a root of a nonzero bivariate polynomial $P(x,y)\in k[x,y]$. It is known that such an algebraic function is also $D$-finite (abeloeuvres, , p. 287), and we call differential resolvent an operator annihilating all the roots of $P$. Cockle’s algorithm (cockle1861transcendental, ; bostan2007algeqtodiffeq, ) computes a differential resolvent as a linear relation between the images of $\alpha^{(i)}$ in the quotient ring $k(x)[y]/(P)$.
Least common left multiple and symmetric product
$D$-finite functions are closed under sums and products, and the algorithms derived from the proof are expressible as instances of Problem 1 (stanley1980dfinite, , Thm. 2.3). Efficient algorithms for these two operations are needed for manipulating $D$-finite functions represented by annihilating operators (salvy2019survey, ; salvy1994gfun, ). For $L_{1},\dots,L_{s}\in k[x]\left<\partial_{x}\right>$, a minimal order linear differential operator that annihilates all the sums (resp. products) of solutions of the $L_{i}$’s is called a least common left multiple (LCLM) (resp. symmetric product).
Our thesis is that studying Problem 1 can contribute to a deeper understanding of the four specific instances mentioned above, since this common formulation underlines a shared structure between these problems. Following this idea, we specifically propose a uniform approach to prove better degree bounds on the polynomials $\eta_{i}$ in (2) (or to retrieve the best known bounds) for these four instances at once. Such bounds are useful as they provide estimates on the output size of algorithms adressing these problems, and thus determine the notion of optimal algorithm in an algebraic model of complexity. In addition, there are examples of algorithms whose running time is dependent on the output size (vanderHoeven2016complexity_skew_arith, ; bostan2007algeqtodiffeq, ). In this setting, proving tight degree bounds is of key importance for complexity analysis.
A straightforward approach to obtain degree bounds is to consider $\eta$ as the solution of the linear system defined in (2) and to derive degree bounds using classical techniques such as Cramer’s rule (bostan2010complexity, ; kauers2014bounds_Dfinite_closure, ). By doing so, if $D$ is a degree bound on both the numerator and the denominator of $T$, one can show that the coefficients of the solution of Problem 1 have degree bounded by $O(n\deg a+n^{2}D)$. The precise result is stated and proved in Section 3.
Experiments suggest that this bound is reached for generic instances of Problem 1. However for our four instances, in which the matrix $T$ possesses a structure, the bound above is an overestimation of the actual degrees. For instance, let us consider the case of the algorithm based on Hermite’s reduction applied to a rational function $f$ with $d_{x}$ (resp. $d_{y}$) a degree bound in $x$ (resp. $y$) on both its numerator and denominator. By applying Cramer’s rule in the linear algebra step of the algorithm, the degree of the telescoper is bounded by $O(d_{y}^{3}d_{x})$ (bostan2010complexity, , Lem. 18). However, an independent analysis based on the Almkvist-Zeilberger algorithm for rational functions (almkvist1990method, ) yields a better bound in $O(d_{y}^{2}d_{x})$ (bostan2010complexity, , Thm. 25). The aim of this work is to explain the gap for each instance we study.
Contributions
We consider representations of a rational matrix $T\in k(x)^{n\times n}$ of the form
(3) $T=W+XM^{-1}Y,$
where $W,X,M,Y$ are polynomial matrices of sizes $n\times n$, $n\times m$, $m\times m$, $m\times n$ respectively and $M$ is non-singular. Such a representation for $T$ is called a realisation (coppel1974matratfun, ) and originates in control theory. We focus here on strictly proper rational matrices: a rational matrix is strictly proper if its limit when $x$ tends to infinity is 00. In the case where $T$ is strictly proper, we have $\det M\cdot T=N$ with $N$ a polynomial matrix of degree $<\deg(\det M)$. We establish a degree bound for the solution of Problem 1, depending on $\deg(\det M)$.
Theorem 1.
Let $T\in k(x)^{n\times n}$ be a strictly proper rational matrix, $\theta=\partial_{x}+T$ and $a\in k[x]^{n}$. Suppose we have a realisation (3) of $T$. Let $\Delta=\det M$, $\delta=\deg\Delta$, and $d_{a}=\deg a$. Then, there exists a solution $\eta=(\eta_{0},\dots,\eta_{\rho})\in k[x]^{\rho+1}\setminus\{0\}$ of Problem 1 for $(T,a)$ satisfying $\deg\eta_{i}\leq\rho d_{a}+\rho\delta-\left(\rho(\rho+1)/2-i\right)$.
The coefficients of the solution of Problem 1 are thus bounded by $O(nd_{a}+n\delta)$. If we choose the trivial realisation $T=(\mathrm{den}\cdot T)(\mathrm{den}\cdot I_{n})^{-1}$ where $\mathrm{den}$ is a least common denominator of the entries of $T$ of degree $D$, we retrieve the previously discussed bound in $O(nd_{a}+n^{2}D)$, since we have $\delta=nD$. In order to prove tight degree bounds, we consider realisations of $T$ minimizing the degree of $\det M$, also known as irreducible realisations (coppel1974matratfun, ). In the four specific instances we focus on, we exhibit a small realisation. Each time, the realisation is obtained by interpreting the map $T$ as carrying out a reduction in a quotient vector space modulo a polynomial or a relation. We prove Theorem 1 in Section 5. It follows from structural properties on the denominators of minors of matrices of rational functions, recalled in Section 4. It is based on a linearisation of the input matrix $T$ also known as a state-space realisation of $T$ (kailath1980linear_systems, ). Besides, the strict properness assumption in Theorem 1 seems to be only required because of the technical tools we use in the proofs. This is at least suggested by our experiments.
In Section 2, Theorem 1 is applied to the four instances mentioned above. For each instance, a degree bound is obtained and compared with the best known bound in the literature. A summary is given in Table 1. For the algorithm based on Hermite’s reduction (row Hermite), the degree bounds depend on $d_{x}$ and $d_{y}$ as defined before. Similarly, for computing the differential resolvent of $P(x,y)$, the degree bounds are given in terms of $d_{x},d_{y}$ the degrees of $P$ in $x$ and $y$ respectively (row AlgeqtoDiffeq). For these two applications, our new bounds are valid under a genericity assumption that is made explicit in Section 2.1. Finally, for LCLM and symmetric products of $s$ linear differential operators of order at most $r$ and degree at most $d$ (rows LCLM and SymProd), the bounds are expressed in terms of the parameters $s,r$ and $d$. The new bounds are established for operators that do not admit the point at infinity as an irregular singularity, i.e. it is either ordinary or a regular singularity (poole1936introduction, , §20). Note that the restrictions on our bounds (genericity, point at infinity not irregular) are only necessary to fulfil the strict properness assumption in Theorem 1. A similar result without this assumption would yield bounds that hold for arbitrary input in our four instances.
Previous Bound New Bound
Hermite $2d_{y}^{2}d_{x}+o(d_{y}^{2}d_{x})$ (bostan2010complexity, ) $2d_{y}^{2}d_{x}+o(d_{y}^{2}d_{x})$  [^1] Under a genericity assumption (see Section 2.1)
AlgeqtoDiffeq $4d_{y}^{2}d_{x}+o(d_{y}^{2}d_{x})$ (bostan2007algeqtodiffeq, ) $2d_{y}^{2}d_{x}+o(d_{y}^{2}d_{x})$ 1
LCLM $ds^{2}r+o(ds^{2}r)$ (bostan2012LCLM, ) $ds^{2}r+o(ds^{2}r)$  [^2] Under the assumption that $x=\infty$ is not an irregular singularity
SymProd $O(dr^{2s})$ (kauers2014bounds_Dfinite_closure, ) $O(dr^{2s-1})$ 2
Table 1. Degree bounds for the four instances
2. Results in specific instances
The proof of our main result is postponed to Section 5. For now we focus on its consequences for proving degree bounds in the problems we have mentioned. In each of the following instances, we exhibit a realisation of the resulting matrix $T$. We then study for which class of inputs the matrix $T$ is strictly proper so that Theorem 1 is applicable.
2.1. Hermite reduction for creative telescoping
Let $f=p/q$ be in $k(x,y)\setminus\{0\}$ with $p$ and $q$ coprime in $k[x,y]$. For simplicity, we assume that $q$ is square-free with respect to $y$, $\deg_{y}p<\deg_{y}q=d_{y}$ and $\deg_{x}p\leq\deg_{x}q=d_{x}$. We also denote $q_{y}=\partial_{y}(q)$ and $q_{x}=\partial_{x}(q)$. The goal is to compute a telescoper $L\in k[x]\left<\partial_{x}\right>\setminus\{0\}$ of $f$, i.e. such that $L(f)=\partial_{y}(h)$ for $h\in k(x,y)$.
2.1.1. Reduction to Problem 1
For any rational function $g\in k(x,y)$ whose denominator is a power of $q$, we denote by $\operatorname{herm}(g)$ its Hermite reduction (hermite1872integration, ), namely the unique polynomial $r(x,y)\in k(x)[y]$ such that $\deg_{y}(r)<\deg_{y}(q)$ and $g=\partial_{y}(h)+r/q$ for $h\in k(x,y)$. Let $\mathcal{A}$ be the quotient $k(x)[y]/(q)$. It is known that $\operatorname{herm}$ defines a $k(x)$-linear map from $k(x)[y,q^{-1}]$ to $\mathcal{A}$ and $\operatorname{herm}(g)=0$ if and only $g$ is a derivative (wrt. $y$). Moreover, one can show (bostan2010complexity, ) that $\operatorname{herm}(\partial_{x}\cdot g)=(\partial_{x}+T)\cdot\operatorname{herm}(g)$ where $T$ is the $k(x)$-linear map:
$T\colon a\in\mathcal{A}\mapsto-\operatorname{herm}(q_{x}a/q^{2})\in\mathcal{A}.$
A telescoper $L$ of $f$ is such that $\operatorname{herm}(L\cdot f)=0$ and thus by linearity of $\operatorname{herm}$ it can be seen as a linear relation between $\operatorname{herm}(\partial_{x}^{i}\cdot f)$ for $i\geq 0$. Besides, by induction on $i\geq 0$, we obtain $\operatorname{herm}(\partial_{x}^{i}\cdot f)=\theta^{i}(p)$ with $p$ the numerator of $f$ and $\theta$ the pseudo-linear map $\partial_{x}+T$. Therefore, a minimal telescoper for $f$ can be obtained as the solution of Problem 1 on input $(T,p)$. This is essentially the algorithm in (bostan2010complexity, , Sec. 3.1).
2.1.2. Realisation of $T$
In the basis $(1,y,\dots,y^{d_{y}-1})$, we consider $T$ as a matrix with rational function as coefficients, i.e. $T\in k(x)^{d_{y}\times d_{y}}$.
Lemma 0.
There exist polynomial matrices $X,M$ and $Y$, with $M$ invertible and $\deg\det M\leq 2d_{x}d_{y}$ such that $T=XM^{-1}Y$.
Proof.
By analysing Hermite reduction for $q_{x}a/q^{2}$ and its associated linear system as in (horowitz1971algorithms, ; bostan2010complexity, ), for all $a\in\mathcal{A}$ we have $q_{x}a/q^{2}=\partial_{y}(A/q)+r/q$ for $A,r\in\mathcal{A}$ and $r=\operatorname{herm}(q_{x}a/q^{2})$. This can be reformulated as
(4) $q_{x}a=q\partial_{y}(A)-q_{y}A+qr,$
just by multiplying the previous equality by $q^{2}$. Viewing $A$ and $r$ as undetermined polynomials in $\mathcal{A}$, $(A,r)$ is solution to the linear system
(5) $\begin{bmatrix}M_{1}&M_{2}\end{bmatrix}\begin{bmatrix}A\\ r\end{bmatrix}=q_{x}a,$
where $M_{1},M_{2}\in k[x]^{2d_{y}\times d_{y}}$ of degree at most $d_{x}$. By (bostan2010complexity, , Lem. 6), $M=\begin{bmatrix}M_{1}&M_{2}\end{bmatrix}$ is invertible and by a direct expansion, its determinant has degree $\leq 2d_{x}d_{y}$. Let $Y$ denote the multiplication by $-q_{x}$, i.e. $Y\colon a\mapsto-q_{x}a$. For all $a\in\mathcal{A}$, we have $Ta=XM^{-1}Ya$ where $X$ is simply the projection according to the last $d_{y}$ coordinates. ∎
2.1.3. Strict properness
It remains to study when the matrix $T$ is strictly proper in order to apply Theorem 1 with the realisation of Lemma 1. Let $\Delta=\det M$. Since $T=XM^{-1}Y$ with $X$ of degree 0 and $Y$ of degree at most $d_{x}-1$, it is sufficient that $M^{-1}=O(1/x^{d_{x}})$ at infinity for $T$ to be strictly proper. However $M^{-1}=1/\Delta\cdot M^{*}$ where $M^{*}$ is the adjugate matrix of $M$. So the entries of $M^{*}$ are minors of $M$ of dimension $12d_{y}-$ 1. Besides, the entries of $M$ have degree at most $d_{x}$, thus $\deg\Delta\leq 2d_{x}d_{y}$ and the entries of $M^{*}$ all have degree at most $(2d_{y}-1)d_{x}$. Hence, it suffices that $\deg\Delta=2d_{x}d_{y}$ to have $M^{-1}=O(1/x^{d_{x}})$ and $T$ strictly proper.
Lemma 0.
Up to a sign, $\Delta=\operatorname{lc}(q)\operatorname{Res}_{y}(q,q_{y})$, where $\operatorname{lc}(q)$ is the leading coefficient of $q$ in $y$ and $\operatorname{Res}_{y}$ denotes the resultant wrt. $y$.
Proof.
Let $\mathcal{B}$ be the vector space $k(x)[y]_{\leq 2d_{y}-1}$. Recall that $M$ is the matrix of the system described in (4) and (5). We have $M=\begin{bmatrix}M_{1}&M_{2}\end{bmatrix}$ with $M_{2}$ the matrix of the multiplication by $q$, i.e. $M_{2}\colon r\in\mathcal{A}\mapsto qr\in\mathcal{B}$, and $M_{1}$ the matrix of $A\in\mathcal{A}\mapsto(q\partial_{y}(A)-q_{y}A)\in\mathcal{B}$. Therefore, one can write $M_{1}=M_{2}\cdot U-M_{3}$ with $M_{3}$ the matrix of the multiplication by $q_{y}$, i.e. $M_{3}\colon A\in\mathcal{A}\mapsto q_{y}A\in\mathcal{B}$, and $U$ the matrix of $A\in\mathcal{A}\mapsto\partial_{y}(A)\in\mathcal{A}$. Writing in the basis $(1,\dots,y^{d_{y}-1})$ of $\mathcal{A}$,
$U={\begin{bmatrix}0&1&\\ &0&2\\ &&\ddots&\ddots\\ \end{bmatrix}}.$
Hence every column of $M_{2}\cdot U$ is a multiple of a column of $M_{2}$. So,
$\Delta=\det\begin{bmatrix}M_{1}&M_{2}\end{bmatrix}=\det\begin{bmatrix}-M_{3}&M_{2}\end{bmatrix}.$
We just proved that, up to a sign, $\Delta$ is the determinant of $(A,r)\in\mathcal{A}^{2}\mapsto q_{y}A+qr\in\mathcal{B}$. The matrix of the previous map in the bases $((1,0),\dots,(y^{d_{y}-1},0),(0,1),\dots,(0,y^{d_{y}-1}))$ and $(1,\dots,y^{2d_{y}-1})$ is
$\begin{bmatrix}S&*\\ 0&\operatorname{lc}(q)\end{bmatrix},$
where $S$ is a matrix representing the Sylvester map associated to $(q_{y},q)$. Finally, up to a sign, $\Delta=\operatorname{lc}(q)\operatorname{Res}_{y}(q,q_{y})$. ∎
As a consequence, $\deg\Delta$ has maximal degree $2d_{x}d_{y}$ if and only if $\deg(\operatorname{lc}(q))=d_{x}$ and $\operatorname{Res}_{y}(q,q_{y})$ has maximal degree $(2d_{y}-1)d_{x}$.
Lemma 0.
Let $q\in k[x,y]$ be square-free with respect to $y$ of degree $d_{x}$ and $d_{y}$ in $x$ and $y$ respectively. The resultant $\operatorname{Res}_{y}(q,q_{y})$ has maximal degree $(2d_{y}-1)d_{x}$ iff the coefficient of degree $d_{x}$ in $x$ of $q$, which is a polynomial in $y$, has degree $d_{y}$ and is square-free.
Proof.
Let $S$ be the Sylvester matrix associated to $(q,q_{y})$. One can write $S=x^{d_{x}}S_{0}+O(x^{d_{x}-1})$ with $S_{0}\in k^{(2d_{y}-1)\times(2d_{y}-1)}$. So $\operatorname{Res}_{y}(q,q_{y})$ has maximal degree $(2d_{y}-1)d_{x}$ if and only if $\det S_{0}\neq 0$. Let $Q\in k[y]$ be the coefficient of $q$ of degree $d_{x}$ in $x$. $S_{0}$ is the matrix of the map $(U,V)\in k[y]_{\leq d_{y}-2}\times k[y]_{\leq d_{y}-1}\mapsto UQ+V\partial_{y}(Q)$. This map is invertible if and only if $Q$ has degree $d_{y}$ and is square-free, whence the result. ∎
In summary, if the coefficient of degree $d_{x}$ in $x$ of $q$ has degree $d_{y}$ and is square-free, then $T$ is strictly proper.
2.1.4. Degree bound
Theorem 4.
Let $f(x,y)=p/q\in k(x,y)$ with $q$ square-free with respect to $y$, $(d_{x},d_{y})=(\deg_{x}(q),\deg_{y}(q))$ and $\deg_{x}p\leq d_{x}$, $\deg_{y}p<d_{y}$. If the coefficient of degree $d_{x}$ in $x$ of $q$ has degree $d_{y}$ and is square-free, then there exists a minimal telescoper for $f$ of order $r\leq d_{y}$ and degree at most $rd_{x}+2rd_{y}d_{x}-1/2\cdot r(r-1)$.
Proof.
Under this assumption, the matrix $T$ is strictly proper and by Lemma 1, there is a realisation $T=XM^{-1}Y$ with $\deg\det M\leq 2d_{x}d_{y}$. Thus by Theorem 1, the claimed degree bound holds. ∎
In Theorem 4, the condition on the coefficient of degree $d_{x}$ in $x$ of $q$ holds when the coefficients of $q$ do not belong to a certain nontrivial hypersurface. In other words, this is a genericity assumption. Thus, in the generic situation, we retrieve the best known bound on the degree of a minimal telescoper of a bivariate rational function (bostan2010complexity, , Thm. 25), namely a degree governed by $2d_{y}^{2}d_{x}$. This bound was derived from the analysis of the Almkvist-Zeilberger algorithm for rational functions (almkvist1990method, ). This is the first time a degree bound in $O(d_{y}^{2}d_{x})$ is derived directly from the Hermite reduction-based algorithm.
2.2. Differential resolvent
Let $P(x,y)\in k[x,y]$ be square-free with respect to $y$ and let $d_{x}$ and $d_{y}$ be the degrees in $x$ and $y$ of $P$. We write $P_{x}=\partial_{x}(P)$ and $P_{y}=\partial_{y}(P)$. In this section, $\mathcal{A}$ denotes the quotient ring $k(x)[y]/(P)$.
2.2.1. Reduction to Problem 1
Following the lines of (bostan2007algeqtodiffeq, , Sec. 2.1), we first show that a differential resolvent of $P$ can be computed as a solution of Problem 1. Let $\alpha_{1},\dots,\alpha_{d_{y}}$ be the roots of $P$ in the algebraic closure of $k(x)$. For any root $\alpha$ of $P$, by differentiating $P(x,\alpha)=0$ and since $P$ is square-free, we have $\partial_{x}(\alpha)=-P_{x}(x,\alpha)/P_{y}(x,\alpha)$. Hence the derivative $\partial_{x}$ uniquely extends to $k(x)(\alpha_{1},\dots,\alpha_{d_{y}})$. By induction, for all $i\geq 1$, there exists a polynomial $C_{i}\in k[x,y]$ such that $\alpha^{(i)}=C_{i}(x,\alpha)/P_{y}(x,\alpha)^{2i-1}$, for any root $\alpha$ of $P$. Since $P$ and $P_{y}$ are coprime, one can define $D_{i}\in\mathcal{A}$ that is the image of $C_{i}/P_{y}^{2i-1}$ modulo $P$. If one defines $D_{0}=y$, the sequence $(D_{i})_{i\geq 0}$ is such that $D_{i}(\alpha)=\alpha^{(i)}$ for all roots $\alpha$ of $P$. And thus it suffices to find the minimal linear dependence between the $D_{i}$’s to compute the differential resolvent of $P$.
Moreover, by differentiating $\alpha^{(i)}=D_{i}(x,\alpha)$, we obtain
$\alpha^{(i+1)}$ $=\partial_{x}(D_{i}(x,\alpha))+\partial_{x}(\alpha)\cdot\partial_{y}(D_{i}(x,\alpha))$
$=\partial_{x}(D_{i}(x,\alpha))-\frac{P_{x}(x,\alpha)}{P_{y}(x,\alpha)}\partial_{y}(D_{i}(x,\alpha)).$
Therefore $D_{i+1}=(\partial_{x}+T)\cdot D_{i}$, where $T$ is the $k(x)$-linear map $T\colon a\mapsto-\partial_{y}(a)P_{x}/P_{y}~{}\mathrm{mod}~{}P$ on $\mathcal{A}$. Finally, the differential resolvent of $P$ is the solution of Problem 1 for $(T,y)$.
2.2.2. Realisation of $T$
Fix the monomial basis $(1,y,\dots,y^{d_{y}-1})$ of $\mathcal{A}$ and view $T$ as a matrix in $k(x)^{d_{y}\times d_{y}}$.
Lemma 0.
There exist polynomial matrices $X,M$ and $Y$, with $M$ invertible and $\deg\det M\leq(2d_{y}-1)d_{x}$ such that $T=XM^{-1}Y$.
Proof.
Recall that $T\cdot a$ is the solution $V$ of the following Bézout equation
$-\partial_{y}(a)P_{x}=U\cdot P+V\cdot P_{y}.$
So if $Y\colon a\in\mathcal{A}\mapsto-\partial_{y}(a)P_{x}\in k(x)[y]_{\leq 2d_{y}-2}$ and $X$ is the projection $(U,V)\mapsto V$ onto the last $d_{y}$ coordinates, we have $T=XM^{-1}Y$ where $M$ is the Sylvester map associated to $P$ and $P_{y}$ and $X,M,Y$ are three polynomial matrices. So $\det M$ is the resultant of $P$ and $P_{y}$ (with respect to $y$) and it has degree at most $(2d_{y}-1)d_{x}$. ∎
2.2.3. Strict properness
An analysis similar to that of Section 2.1 shows that it suffices that $M^{-1}=O(1/x^{d_{x}})$ for $T$ to be strictly proper. This is in particular the case when $\det M$ has maximal degree $(2d_{y}-1)d_{x}$, i.e., when $\operatorname{Res}_{y}(P,P_{y})$ has maximal degree. By Lemma 3, this property holds when the coefficient of degree $d_{x}$ of $P$ in $x$ has degree $d_{y}$ and is square-free.
2.2.4. Degree bound
Theorem 6.
Let $P\in k[x,y]$ be square-free with respect to $y$, $(d_{x},d_{y})=(\deg_{x}P,\deg_{y}P)$. If the coefficient of degree $d_{x}$ of $P$ in $x$ has degree $d_{y}$ and is square-free, then there exists a differential resolvent of $P$ of order $r\leq d_{y}$ and degree at most $r(2d_{y}-1)d_{x}-1/2\cdot r(r-1)$.
Proof.
Under this condition, the matrix $T$ is strictly proper. Moreover, it has realisation $T=XM^{-1}Y$ with $\deg\det M\leq(2d_{y}-1)d_{x}$ by Lemma 5. The claimed bound follows from Theorem 1. ∎
Again, the condition above is a genericity assumption on $P$. Our bound is in $O(d_{y}^{2}d_{x})$ and governed by $2d_{y}^{2}d_{x}$. The previous best known bound behaves like $4d_{y}^{2}d_{x}$ (bostan2007algeqtodiffeq, , Thm. 1) and its proof is independent from Cockle’s algorithm. So this is an improvement of the constant from $4$ to $2$ in the generic situation.
Experimentally if we assume that $d_{x}=d_{y}=d$, we observe that the output degree is bounded by $d(2d^{2}-3d+3)$. Our bound gives $d(2d^{2}-3d/2+1/2)$ and the previous bound is $d(4d^{2}-11d/2+7/2)$ (bostan2007algeqtodiffeq, ). This is the first time that a bound with the first term $2d^{3}$ is obtained.
2.3. Least common left multiple
Let $L_{1},\dots,L_{s}$ be linear differential operators in $k[x]\left<\partial_{x}\right>$ and let $L=\operatorname{LCLM}(L_{1},\dots,L_{s})$ be a least common left multiple. Again, we apply Theorem 1 to bound the degree of $L$ and compare with the existing bound (bostan2012LCLM, ). For simplicity, we first assume $s=2$ and $L=\operatorname{LCLM}(L_{1},L_{2})$ as the approach will be easily generalised for arbitrary $s$.
2.3.1. Reduction to Problem 1 and realisation
Let $r_{1}$, $r_{2}$ be the respective orders of $L_{1}$ and $L_{2}$, and $\alpha=\alpha_{1}+\alpha_{2}$ be the sum of a solution of $L_{1}$ and $L_{2}$. The coefficients of $L$ are read from a linear relation between the successive derivatives of $\alpha$. These derivatives all lie in the finite dimensional vector space spanned by $A=(\alpha_{1},\dots\alpha_{1}^{(r_{1}-1)},\alpha_{2},\dots,\alpha_{2}^{(r_{2}-1)})$. Let $L_{1}=p_{1,r_{1}}\partial_{x}^{r_{1}}+\cdots+p_{1,0}$ and $L_{2}=p_{2,r_{2}}\partial_{x}^{r_{2}}+\cdots+p_{2,0}$ with $p_{1,r_{1}}$ and $p_{2,r_{2}}$ nonzero. By induction suppose that for $\ell\geq 0$, we have written $\alpha^{(\ell)}=\sum_{j=0}^{r_{1}-1}a_{j}\alpha_{1}^{(j)}+\sum_{j=0}^{r_{2}-1}b_{j}\alpha_{2}^{(j)}$ with $a_{j},b_{j}\in k(x)$. Then differentiation yields
$\alpha^{(\ell+1)}=\sum_{j=0}^{r_{1}-1}\left(a_{j}^{\prime}\alpha_{1}^{(j)}+a_{j}\alpha_{1}^{(j+1)}\right)+\sum_{j=0}^{r_{2}-1}\left(b_{j}^{\prime}\alpha_{2}^{(j)}+b_{j}\alpha_{2}^{(j+1)}\right).$
Finally, we use $L_{1}(\alpha_{1})=L_{2}(\alpha_{2})=0$ to rewrite $\alpha_{1}^{(r_{1})}$ and $\alpha_{2}^{(r_{2})}$ on the generating set $A$. So if $V_{\ell}$ denotes the vector of coefficients of $\alpha^{(\ell)}$ in $A$ for all $\ell$, we obtain $V_{\ell+1}=(\partial_{x}+T)\cdot V_{\ell}$, where $T\in k(x)^{(r_{1}+r_{2})\times(r_{1}+r_{2})}$ is the block diagonal matrix $\operatorname{Diag}(C_{1},C_{2})$ with $C_{i}$ the companion matrix associated to $L_{i}$, i.e. $C_{i}\in k(x)^{r_{i}\times r_{i}}$ is companion with last column $-1/p_{i,r_{i}}\cdot\begin{bmatrix}p_{i,0}&\cdots&p_{i,r_{i}-1}\end{bmatrix}^{t}$. But such a block companion matrix $T$ is clearly not strictly proper since some of its entries are $1$.
In order to be able to apply Theorem 1 we propose to express the problem in a different generating set. Let $\delta=x\partial_{x}$, also known as the Euler operator. We choose to write the successive derivatives of $\alpha$ as linear combinations of elements of the new generating set $B=(\alpha_{1},\dots,\delta^{r_{1}-1}\alpha_{1},\alpha_{2},\dots,\delta^{r_{2}-1}\alpha_{2})$. Now if $W_{\ell}$ denotes the vector of coefficients of $\alpha^{(\ell)}$ in $B$ for all $\ell$, by differentiating we get that $W_{\ell+1}=(\partial_{x}+T_{\delta})\cdot W_{\ell}$ with $T_{\delta}\in k(x)^{(r_{1}+r_{2})\times(r_{1}+r_{2})}$ the matrix whose columns are the derivatives of the elements of $B$ expressed on $B$. Notice that for $i=1,2$ and $\ell\geq 0$, we have $(\delta^{\ell}\alpha_{i})^{\prime}=1/x\cdot\delta^{\ell+1}\alpha_{i}$. Also, any linear differential operator in $x$ and $\partial_{x}$ can be rewritten as a linear differential operator of the same order in $x$ and $\delta$ with the following degree bounds on the coefficients.
Lemma 0.
Let $L=\sum_{j=0}^{r}p_{j}(x)\partial_{x}^{j}\in k[x]\left<\partial_{x}\right>$ of order $r$ with $d_{j}=\deg p_{j}$. There exists an operator $\tilde{L}=\sum_{j=0}^{r}q_{j}(x)\delta^{j}\in k[x]\left<\delta\right>$ of order $r$ whose solution set is the same as $L$. Moreover, we have $\deg q_{r}=d_{r}$ and $\deg q_{j}\leq r+\max_{\ell\geq j}(d_{\ell}-\ell)$ for all $j$.
Proof.
The result is a direct consequence of the well-known identity $x^{j}\partial_{x}^{j}=(\delta-j+1)\cdots\delta$ for all $j\geq 0$ (poole1936introduction, , p. 26). ∎
By Lemma 7, $L_{i}=q_{i,r_{i}}(x)\delta^{r_{i}}+\cdots+q_{i,0}(x)$ with $q_{i,r_{i}}$ nonzero for $i=1,2$. Let $D_{i}$ be the companion matrix of size $r_{i}$ whose last column is $-1/q_{i,r_{i}}\cdot\begin{bmatrix}q_{i,0}&\cdots&q_{i,r_{i}-1}\end{bmatrix}^{t}$. One can write $T_{\delta}=1/x\cdot\operatorname{Diag}(D_{1},D_{2})$. Therefore $T_{\delta}=XM^{-1}$ with
$M=x\cdot\operatorname{Diag}(1,\dots,1,-q_{1,r_{1}},1,\dots,1,-q_{2,r_{2}}),$
and $X$ the block diagonal matrix whose blocks are companion matrices associated to $(q_{1,0},\dots,q_{1,r_{1}-1})$ and $(q_{2,0},\dots,q_{2,r_{2}-1})$.
2.3.2. Strict properness
The matrix $T_{\delta}$ is strictly proper when $\operatorname{Diag}(D_{1},D_{2})=O(1)$ and thus when $\deg(q_{i,j})\leq\deg(q_{i,r_{i}})$ for all $j$ and $i=1,2$. By Lemma 7, we deduce that $\deg(q_{i,j})\leq\deg(q_{i,r_{i}})$ when $\deg(p_{i,j})+r_{i}-j\leq\deg(p_{i,r_{i}})$ for all $j$. So if $d$ is the degree of $L_{i}$, we need $\deg(p_{i,j})\leq d-(r_{i}-j)$. This is equivalent to $x=\infty$ not being an irregular singularity of the equation defined by $L_{i}$ (poole1936introduction, , §20). Hence, we conclude that if $x=\infty$ is not an irregular singularity of $L_{1}$ nor of $L_{2}$, then $T_{\delta}$ is strictly proper.
2.3.3. Degree bound
Theorem 8.
Let $L_{1},L_{2}\in k[x]\left<\partial_{x}\right>$ be of respective order $r_{1}$ and $r_{2}$, and degree at most $d$ in $x$. Suppose $x=\infty$ is not an irregular singularity of $L_{1}$ nor of $L_{2}$. Then there exists a LCLM $L$ of $L_{1}$ and $L_{2}$ of order $r\leq r_{1}+r_{2}$ and degree at most $r(r_{1}+r_{2}+2d)-1/2\cdot r(r-1)$.
Proof.
The coefficients of $L$ can be directly derived from the solution to Problem 1 with input $(T_{\delta},a)$ where
$a=(\overbrace{1,0,\dots,0}^{r_{1}},\overbrace{1,0,\dots,0}^{r_{2}})^{t}.$
Besides, $T_{\delta}$ is strictly proper by assumption. Also, $T_{\delta}=XM^{-1}$ with $\deg\det M\leq 2d+r_{1}+r_{2}$. Thus the result follows from Theorem 1. ∎
This result generalizes with $s\geq 2$ differential equations with no irregular singularity at infinity and a matrix $T_{\delta}$ with $s$ different blocks. Let $r_{i}$ denote the order of $L_{i}$ for $1\leq i\leq s$, $R=\sum_{i=1}^{s}r_{i}$ and $d$ be a bound on the degrees in $x$ of all the $L_{i}$’s. A least common left multiple $L=\operatorname{LCLM}(L_{1},\dots,L_{s})$ can be computed as the solution of Problem 1 for $T_{\delta}=XM^{-1}\in k(x)^{R\times R}$ and
$a=(\overbrace{1,0,\dots,0}^{r_{1}},\dots,\overbrace{1,0,\dots,0}^{r_{s}})^{t},$
where $X,M$ are polynomial matrices. This time $\Delta=\det M$ is the product of $x^{R}$ and the leading coefficients of the $L_{i}$’s (up to a sign), so $\deg\Delta\leq sd+R$. Hence by Theorem 1, the order of $L$ is less than $R$ and its degree in $x$ is bounded by $R(sd+R)$. This is slightly larger than the bound $d(s(R+1)-R)$ (bostan2012LCLM, , Thm. 6), but as $R\leq sd$ since $x=\infty$ is not an irregular point, the leading term of both bounds behaves like $dsR\leq ds^{2}\max_{i}(r_{i})$.
2.4. Symmetric product
Let $L_{1},L_{2}\in k[x]\left<\partial_{x}\right>$ and $L=L_{1}\otimes L_{2}$ be a symmetric product. Let $r_{i}$ be the order of $L_{i}$ and $d_{i}$ be its degree for $i=1,2$. For $i=1,2$, we write $L_{i}=\sum_{j=0}^{r_{i}}p_{i,j}\partial_{x}^{j}$ with $p_{i,j}\in k[x]$. It is known that the order of $L$ is at most $r_{1}r_{2}$ and the degree of $L$ in $x$ can be bounded by $r_{1}^{2}r_{2}^{2}(d_{1}+d_{2})$ (kauers2014bounds_Dfinite_closure, , Thm. 8). However, experiments suggest that this bound is not tight since one expects the degree of $L$ to be bounded by $(r_{1}r_{2}-r_{1}-r_{2}+2)(d_{1}r_{2}+d_{2}r_{1})$.
2.4.1. Reduction to Problem 1
By Leibniz’s rule, we express the successive derivatives of $\alpha=\alpha_{1}\alpha_{2}$ according to the elements $(\delta^{h}\alpha_{1}\delta^{p}\alpha_{2})$ for $0\leq h<r_{1}$ and $0\leq p<r_{2}$. We again use the Euler operator to satisfy the strict properness assumption. Let $b_{h,p}=\delta^{h}\alpha_{1}\delta^{p}\alpha_{2}$ for all $h,p$ and suppose that we have
$\alpha^{(\ell)}=\sum_{h=0}^{r_{1}-1}\sum_{p=0}^{r_{2}-1}e_{h,p}(x)b_{h,p},$
with $e_{h,p}\in k(x)$. We denote by $V_{\ell}\in k(x)^{r_{1}r_{2}}$ the vector of coefficients of $\alpha^{(\ell)}$ in this decomposition. Differentiating gives
$\alpha^{(\ell+1)}=\sum_{h,p}\left(e_{h,p}^{\prime}b_{h,p}+1/x\cdot e_{h,p}(b_{h+1,p}+b_{h,p+1})\right).$
Next the relations $L_{1}(\alpha_{1})=L_{2}(\alpha_{2})=0$ are used to rewrite $b_{r_{1},p}$ and $b_{h,r_{2}}$ in the expression above. Thus, one can write $V_{\ell+1}=(\partial_{x}+T)\cdot V_{\ell}$ where $T$ is $k(x)$-linear and maps any $b_{h,p}$ for $0\leq h<r_{1}$ and $0\leq p<r_{2}$ to $1/x\cdot(b_{h+1,p}+b_{h,p+1})$ rewritten according to $(b_{h,p})$ for $0\leq h<r_{1}$ and $0\leq p<r_{2}$ using $L_{1}$ and $L_{2}$. For $i=1,2$, let $L_{i}=\sum_{j=0}^{r_{i}}q_{i,j}\delta^{i}$ with $q_{i,j}\in k[x]$. In the basis $(b_{0,0},\dots,b_{0,r_{2}-1},b_{1,0},\dots,b_{r_{1}-1,r_{2}-1})$, $T$ can be seen as the matrix
(6) $T=1/x\cdot(D_{1}\otimes I_{r_{2}}+I_{r_{1}}\otimes D_{2})\in k(x)^{r_{1}r_{2}\times r_{1}r_{2}},$
where $D_{i}$ denotes the same companion matrix as in Section 2.3 and $\otimes$ the Kronecker product. Also, the coefficients of $L$ can be directly read from the solution to Problem 1 with input $(T,e_{1})$ where $e_{1}=(1,0,\dots)\in k(x)^{r_{1}r_{2}}$. Besides, if the order of $L$ is $r_{1}r_{2}$, we retrieve Amitsur’s notion of resultant of the differential polynomials $L_{1}$ and $L_{2}$ (amitsur1954diffpoly_divalgebra, , § 4).
2.4.2. Realisation of $T$
We have $T=(1/x\cdot D_{1})\otimes I_{r_{2}}+I_{r_{1}}\otimes(1/x\cdot D_{2})$. By Section 2.3.1, $1/x\cdot D_{i}$ can be written as $X_{i}M_{i}^{-1}$ with $X_{i}\in k[x]^{r_{i}\times r_{i}}$ the companion matrix whose last column is $\begin{bmatrix}q_{i,0}&\cdots&q_{i,r_{i}-1}\end{bmatrix}^{t}$ and $M_{i}=\operatorname{Diag}(x,\dots,x,-q_{i,r_{i}}x)\in k[x]^{r_{i}\times r_{i}}$. Moreover, we have $\det M_{i}=-x^{r_{i}}q_{i,r_{i}}$. Next, $(1/x\cdot D_{1})\otimes I_{r_{2}}=(X_{1}M_{1}^{-1})\otimes I_{r_{2}}=(X_{1}\otimes I_{r_{2}})\cdot(M_{1}\otimes I_{r_{2}})^{-1}$ and similarly, $I_{r_{1}}\otimes(1/x\cdot D_{2})=(I_{r_{1}}\otimes X_{2})\cdot(I_{r_{1}}\otimes M_{2})^{-1}$. Now let $X=\begin{bmatrix}X_{1}\otimes I_{r_{2}}&I_{r_{1}}\otimes X_{2}\end{bmatrix}$, $M=\operatorname{Diag}(M_{1}\otimes I_{r_{2}},I_{r_{1}}\otimes M_{2})$ and $Y=\begin{bmatrix}I_{r_{1}r_{2}}&I_{r_{1}r_{2}}\end{bmatrix}^{t}$. One can check that $T=XM^{-1}Y$, $\Delta=\det M=(\det M_{1})^{r_{2}}(\det M_{2})^{r_{1}}$ and $\deg\Delta=2r_{1}r_{2}+d_{1}r_{2}+d_{2}r_{1}$.
2.4.3. Strict properness
From (6), it suffices that $D_{i}=O(1)$ for $i=1,2$ for $T$ to be strictly proper. We conclude as in Section 2.3.2: $T$ is strictly proper when both $L_{1}$ and $L_{2}$ do not admit $x=\infty$ as an irregular singularity.
2.4.4. Degree bound
Theorem 9.
Let $L_{1},L_{2}\in k[x]\left<\partial_{x}\right>$ be of respective order $r_{1}$ and $r_{2}$ and degree $d_{1}$ and $d_{2}$. Suppose $x=\infty$ is not an irregular singularity of $L_{1}$ nor of $L_{2}$. Then there exists a symmetric product $L=L_{1}\otimes L_{2}$ of order $r\leq r_{1}r_{2}$ and degree at most $r(2r_{1}r_{2}+d_{1}r_{2}+d_{2}r_{1})-1/2\cdot r(r-1)$.
Proof.
Under this assumption, $T$ is strictly proper. The result is implied by Theorem 1 with $\deg\Delta=2r_{1}r_{2}+d_{1}r_{2}+d_{2}r_{1}$. ∎
Our new bound, although restricted to the assumption that the point at infinity is not irregular, leads to an improvement of the bound in (kauers2014bounds_Dfinite_closure, ) by one order of magnitude. Recall that $d_{1}\geq r_{1}$ and $d_{2}\geq r_{2}$ because of our assumption at infinity, so our bound is in $O(d_{1}^{2}d_{2}^{2})$. This matches the asymptotic behavior we observe in experiments. By contrast, the bound in (kauers2014bounds_Dfinite_closure, ) behaves like $O(d_{1}^{2}d_{2}^{2}(d_{1}+d_{2}))$ for this class of inputs.
Our approach also generalizes to the computation of a symmetric product of $s\geq 2$ operators satisfying the same condition at infinity. Let $L=L_{1}\otimes\cdots\otimes L_{s}$ with $L_{i}$ of order $r_{i}$ and degree at most $d$ for all $i$. Let $R=r_{1}\cdots r_{s}$ and $r\geq r_{i}$ for all $i$. By assumption, we have $r\leq d$. Let $\alpha=\alpha_{1}\cdots\alpha_{s}$ where $L_{i}(\alpha_{i})=0$ for all $i$. By the Leibniz rule, one can express the successive derivatives of $\alpha$ as linear combinations of $(\alpha_{1}^{(h_{1})}\cdots\alpha_{s}^{(h_{s})})$ for $0\leq h_{i}<r_{i}$, $1\leq i\leq s$. Therefore $L$ can still be derived from a solution of Problem 1 with $T\in k(x)^{R\times R}$. And one can write $T=XM^{-1}Y$ generalizing the proof of Theorem 9, with $\deg\det M\leq\sum_{i=1}^{s}(r_{1}\cdots r_{i-1}(r_{i}+d)r_{i+1}\cdots r_{s})$. Then, $\deg\det M\leq s(R+dr^{s-1})$ so $L$ is of order at most $R$ and degree bounded by $sR(R+dr^{s-1})=O(dr^{2s-1})$. One should compare with the previous best known bound $sR^{2}d=O(dr^{2s})$ (kauers2014bounds_Dfinite_closure, , Thm. 8).
3. Direct bound by Cramer’s rule
For the purpose of explaining our approach to show Theorem 1, we start by proving a bound that is directly obtained from the formulation of Problem 1: $\eta$ is a solution of a linear system. This is the bound one should expect by adopting the most natural viewpoint on the input rational matrix $T\in k(x)^{n\times n}$, i.e. viewing $T$ as its numerator $N$ which is a polynomial matrix divided by its denominator $\mathrm{den}\in k[x]$.
Theorem 1.
Let $T\in k(x)^{n\times n}$ with denominator $\mathrm{den}$, and $a\in k[x]^{n}$. Let $d=\deg(\mathrm{den})$, $D=\deg(\mathrm{den}\cdot T)$, $d_{a}=\deg a$ and $\tilde{D}=\max(d-1,D)$. Then, there exists a solution $\eta=(\eta_{0},\dots,\eta_{\rho})\in k[x]^{\rho+1}\setminus\{0\}$ of Problem 1 for $(T,a)$ of the form $\eta_{i}=\mathrm{den}^{i}p_{i}$ with $p_{i}\in k[x]$ of degree at most $\rho d_{a}+(1/2\cdot\rho(\rho+1)-i)\tilde{D}$.
Proof.
By induction on $i\geq 0$, we show that $\theta^{i}a=b_{i}/\mathrm{den}^{i}$ with $b_{i}\in k[x]^{n}$ of degree at most $d_{a}+i\tilde{D}$. This is true for $i=0$ and $\theta^{i+1}a=\partial_{x}(b_{i}/\mathrm{den}^{i})+T\cdot b_{i}/\mathrm{den}^{i}=1/\mathrm{den}^{i+1}(\mathrm{den}\cdot b_{i}^{\prime}-i\mathrm{den}^{\prime}\cdot b_{i}+(\mathrm{den}\cdot T)b_{i})$. So we have $b_{i+1}=\mathrm{den}\cdot b_{i}^{\prime}-i\mathrm{den}^{\prime}\cdot b_{i}+(\mathrm{den}\cdot T)b_{i}$, and $\deg(b_{i+1})\leq\deg(b_{i})+\tilde{D}$, thus the claim holds.
Next consider the solution $\nu\in k(x)^{\rho}$ of the linear system $K\nu=-\theta^{\rho}a$ where $K=\begin{bmatrix}a&\theta a&\cdots&\theta^{\rho-1}a\end{bmatrix}\in k(x)^{n\times\rho}$. The solution $\eta$ of Problem 1 can be read from $\nu$ just by multiplying by the denominator of $\nu$. By definition of $\rho$, $\operatorname{rk}(K)=\rho$ so after a deletion of $(n-\rho)$ appropriate rows in the previous system, we obtain an equivalent square system $\tilde{K}\nu=c$, where $\tilde{K}$ and $c$ are obtained from $K$ and $-\theta^{\rho}a$ after the deletion of rows.
By Cramer’s rule, we have $\nu_{i}=\det K_{i}/\det\tilde{K}$ for all $0\leq i<\rho$ where $K_{i}$ is obtained from $\tilde{K}$ after replacing column $i+1$ by $c$. Moreover by a direct expansion we have $\det\tilde{K}=p/\mathrm{den}^{\rho(\rho-1)/2}$ with $p\in k[x]$ of degree at most $\rho d_{a}+1/2\cdot\rho(\rho-1)\tilde{D}$ and $\det K_{i}=p_{i}/\mathrm{den}^{\rho(\rho+1)/2-i}$ with $p_{i}\in k[x]$ of degree at most $\rho d_{a}+(1/2\cdot\rho(\rho+1)-i)\tilde{D}$. So finally for $0\leq i<\rho$, $\nu_{i}=p_{i}/(\mathrm{den}^{\rho-i}p)$.
And thus $\eta_{i}=\mathrm{den}^{\rho}p\cdot\nu_{i}=\mathrm{den}^{i}p_{i}$ for $0\leq i<\rho$ and $\eta_{\rho}=\mathrm{den}^{\rho}p$, whence the result. ∎
In summary, the above proof (which is similar to the approach in (bostan2013cyclic, , §2)) boils down to the study of some of the minors (up to sign) of a matrix $K^{*}$ whose columns are iterates of $\theta$,
$K^{*}=\begin{bmatrix}a&\cdots&\theta^{\rho-1}a&\theta^{\rho}a\end{bmatrix}.$
And we just bounded the size of each minor by simply expanding the determinant and bounding the degrees in every column. For proving Theorem 1, we first show that if the input matrix $T$ has a realisation $T=W+XM^{-1}Y$ then the minors of the matrix $K^{*}$ can always be written with denominator dividing $(\det M)^{\rho}$ (see Proposition 1). Now let $D$ be a degree bound on the denominator of $T$. If we have a small realisation of $T$ meaning that $\deg\det M=O(D)$ then all these minors are actually smaller than naive expansions suggest.
4. Matrices of rational functions
We follow the lines of (coppel1974matratfun, ; kailath1980linear_systems, ) for the main definitions and results we need here.
4.1. Linearisation of matrix fraction description
Let $T\in k(x)^{n\times n}$ with a realisation (3). Such a realisation for $T$ is not unique. For example, for any non-singular $m\times m$ polynomial matrices $D_{1}$ and $D_{2}$, one has the realisation
(7) $T=W+XD_{2}(D_{1}MD_{2})^{-1}D_{1}Y.$
Recall that a matrix $U\in k[x]^{n\times n}$ is unimodular if its determinant is a nonzero element in $k$. When $D_{1}$ or $D_{2}$ are not unimodular the realisation (7) can be reduced.
Let $A,B$ be $m\times p$ and $m\times h$ polynomial matrices. A matrix $D\in k[x]^{m\times m}$ is a common left divisor of $A$ and $B$ if there exist polynomial matrices $A_{1},B_{1}$ such that $A=DA_{1}$ and $B=DB_{1}$. The matrices $A$ and $B$ are said to be left relatively prime if their only left common divisors are unimodular. Similarly, one can define common right divisors and relatively right prime matrices simply by taking transposes in the above definitions. Now a realisation (3) of $T$ is said to be irreducible if $M,Y$ are left relatively prime and $M,X$ are right relatively prime. One can always reduce a given realisation to an irreducible one by the following.
Proposition 0.
(coppel1974matratfun, , Thm. 7-9) If $T$ has the realisation (3) then there exist polynomial matrices $X_{0},M_{0},Y_{0}$ of size $n\times m,m\times m,m\times n$ respectively such that $T$ has the following irreducible realisation
(8) $T=W+X_{0}M_{0}^{-1}Y_{0}.$
Moreover there exist $N,D\in k[x]^{n\times n}$ such that
(9) $T=ND^{-1}$
is an irreducible realisation ($N$ and $D$ are right relatively prime).
A realisation of the form (9) is also called an irreducible (right) matrix fraction description (MFD) for $T$. The proof of Proposition 1 is effective, in the sense that irreducible realisations (8) or (9) can be computed from an arbitrary realisation for $T$.
Now given an MFD (9) for $T$, one can linearise it into a simple form called a state-space realisation (kailath1980linear_systems, ).
Proposition 0.
(kailath1980linear_systems, , Sec. 6.4) Given a right MFD: $T=ND^{-1}\in k(x)^{n\times n}$, let $m=\deg\det D$, one can always construct a so-called state-space realisation
(10) $T(x)=W(x)+B(xI-A)^{-1}C,$
with $W\in k[x]^{n\times n}$ and $B,A,C$ matrices with entries in $k$ of size $n\times m,m\times m,m\times n$.
A state-space realisation (10) is a particular example of a general realisation (3). When $T$ is strictly proper, one has $W=0$ in (10) and the idea of linearisation comes from the fact that we only have a polynomial matrix of degree 1 involved. We use a state-space realisation of $T$ in our proof since this form is convenient when dealing with differentiation. In particular, if $T$ is strictly proper, it has a simple derivative $T^{\prime}=-B(xI-A)^{-2}C$.
4.2. Determinantal denominators
We now recall results about the denominator of minors of rational matrices. For $R\in k(x)^{m\times p}$ a rational matrix and a positive integer $\ell$, we denote by $\varphi_{\ell}(R)$ the monic least common denominator of all minors of $R$ of size at most $\ell$. We also set $\varphi_{0}(R)=1$. The polynomials $\varphi_{\ell}(R)$ are called the determinantal denominators of $R$ (coppel1974matratfun, ). Note that $\varphi_{1}(R)$ is the monic least common denominator of the entries of $R$. So one can write $R=1/\varphi_{1}(R)\cdot N$ with $N$ a polynomial matrix. By definition, for $\ell\geq 0$, $\varphi_{\ell}(R)$ divides $\varphi_{\ell+1}(R)$ and $\varphi_{\ell}(R)=\varphi_{\ell+1}(R)$ for $\ell\geq\operatorname{rk}(R)$. Moreover by a direct expansion of the determinant, $\varphi_{\ell}(R)$ divides $\varphi_{1}(R)^{\ell}$. We recall how determinantal denominators behave with respect to sums, products and inverse of matrices.
Proposition 0.
(coppel1974matratfun, , Thm. 1-2) If a rational matrix $R$ is the sum $R_{1}+R_{2}$ of two rational matrices $R_{1},R_{2}$, then for all $\ell$, $\varphi_{\ell}(R)$ divides $\varphi_{\ell}(R_{1})\varphi_{\ell}(R_{2})$. Moreover, if $\varphi_{1}(R_{1})$ and $\varphi_{1}(R_{2})$ are coprime then equality holds. Similarly if $R$ is the product $R_{1}R_{2}$ of two rational matrices $R_{1},R_{2}$ then for all $\ell$, $\varphi_{\ell}(R)$ divides $\varphi_{\ell}(R_{1})\varphi_{\ell}(R_{2})$.
Proposition 0.
(coppel1974matratfun, , Thm. 4) Let $R$ be a non-singular $m\times m$ rational matrix. If $\det R=c\cdot\alpha/\beta$ with $c\in k^{*}$, $\alpha,\beta\in k[x]$ monic, then $\beta\varphi_{m}(R^{-1})=\alpha\varphi_{m}(R)$.
The case of a realisation (3) is important in this work.
Proposition 0.
(coppel1974matratfun, , Thm. 10) Let $T=W+XM^{-1}Y$ be a realisation of the form (3). For all $\ell\geq 0$, $\varphi_{\ell}(T)$ divides $\varphi_{\ell}(M^{-1})$. If the realisation is irreducible then equality holds.
Corollary 0.
Let $T=W+XM^{-1}Y$ be a realisation of the form (3). For all $\ell\geq 0$, $\varphi_{\ell}(T)$ divides $\det M$.
Proof.
By Proposition 5, $\varphi_{\ell}(T)$ divides $\varphi_{\ell}(M^{-1})$, which divides $\varphi_{m}(M^{-1})$ as $M$ is an $m\times m$ matrix. Next, by Proposition 4, $\varphi_{m}(M^{-1})$ is equal to $\det M\cdot\varphi_{m}(M)$ up to a nonzero constant and $\varphi_{m}(M)=1$ since $M$ is a polynomial matrix. ∎
Proposition 5 and Corollary 6 also apply for MFD (9) or state-space realisation (10) as they are also examples of realisation (3).
5. Degree bound
In this section, we prove our main result (Theorem 1). Let $T\in k(x)^{n\times n}$ be a strictly proper matrix with realisation (3). We define the operator $\theta=\partial_{x}+T$ and take $a\in k[x]^{n}$. We aim at bounding the degrees in the coefficients of the solution of Problem 1 for $(T,a)$. Let $\rho=\dim(\operatorname{span}_{k(x)}(\theta^{i}a,i\geq 0))$ and $\Delta=\det M$.
By Propositions 1 and 2 and the strict properness assumption, one can write an irreducible state-space realisation (10) for $T$ in the form
(10) $T=B(xI-A)^{-1}C\eqcolon BL^{-1}C,$
with $A,B,C$ scalar matrices. Since the above state-space realisation is irreducible, we have $\varphi_{\ell}(T)=\varphi_{\ell}(L^{-1})$ for all $\ell\geq 0$ by Proposition 5 and this divides $\Delta$ by Corollary 6. The degree bound is a consequence of the following result.
Proposition 0.
Let $s_{1}\leq\cdots\leq s_{r}$ be non-negative integers and consider the following matrix of iterates of $\theta$:
(11) $K=\begin{bmatrix}\theta^{s_{1}}a&\cdots&\theta^{s_{r}}a\end{bmatrix}\in k(x)^{n\times r}.$
Then, $\varphi_{\ell}(K)$ divides $\Delta^{s_{r}}$ for all $\ell\geq 0$.
Note that Proposition 1 is somehow unexpected. Indeed, if one expands an arbitrary minor of $K$ of order $r$ by taking $\Delta$ as denominator of $T$, one expects this minor to have denominator $\Delta^{s_{1}+\cdots+s_{r}}$. It is not a priori clear why there should be such simplifications in the denominator of the minors.
5.1. Proof of Proposition 1
Let $\mathcal{O}=k(x)\left<\partial_{x}\right>$ be the ring of linear differential polynomials with rational function coefficients. Then, $\theta=I*\partial_{x}+T$, where $I$ is the identity matrix of order $n$, can be seen as a $n\times n$ matrix with coefficients in $\mathcal{O}$ (jacobson1937pseudolineartransformations, ). Multiplication of matrices of operators, denoted by *, is defined in the usual way by
$\partial_{x}*R=R*\partial_{x}+R^{\prime},$
for any rational matrix $R\in k(x)^{n\times n}$. Next for any matrix $U\in k(x)^{n\times p}$ with columns $u_{1},\dots,u_{p}$, let $\theta(U)$ be the result of applying $\theta$ to each column of $U$, namely
$\theta(U)=\begin{bmatrix}\theta u_{1}&\cdots&\theta u_{p}\end{bmatrix}.$
A simple reasoning by induction proves the following.
Lemma 0.
For all $s\geq 0$, $\theta^{s}=\sum_{j=0}^{s}\binom{s}{j}\theta^{j}(I)*\partial_{x}^{s-j}\in\mathcal{O}^{n\times n}$.
We now compute a convenient expression for the matrices $\theta^{j}(I)$. For $1\leq i\leq j$, we denote by $\Lambda_{i}^{j}$ the product
$\Lambda^{j}_{i}=\prod_{\ell=i}^{j-1}\left[(CB-\ell I)L^{-1}\right],$
with $L=xI-A$ and $A,B,C$ the scalar matrices from (10). We adopt the convention that $\Lambda_{j}^{j}=I$.
Lemma 0.
For $1\leq i\leq j$, $\Lambda^{j+1}_{i}-\left(\Lambda^{j}_{i}\right)^{\prime}=(CB-iI)L^{-1}\Lambda^{j}_{i}$.
Proof.
Let $j\geq 1$, the result holds for $i=j$ since $\Lambda^{j+1}_{j}=(CB-jI)L^{-1}$ and $\Lambda_{j}^{j}=I$. Suppose it holds for $1<i+1\leq j$ and recall that $(L^{-1})^{\prime}=-L^{-2}$. At rank $i$, we obtain
$\Lambda^{j+1}_{i}-\left(\Lambda^{j}_{i}\right)^{\prime}=(CB-iI)\left[L^{-1}\Lambda^{j+1}_{i+1}-(L^{-1}\Lambda^{j}_{i+1})^{\prime}\right]$
$=(CB-iI)L^{-1}\left[\Lambda^{j+1}_{i+1}-(\Lambda^{j}_{i+1})^{\prime}+L^{-1}\Lambda^{j}_{i+1}\right].$
And by the induction hypothesis, the following concludes the proof,
$\Lambda^{j+1}_{i+1}-(\Lambda^{j}_{i+1})^{\prime}+L^{-1}\Lambda^{j}_{i+1}=(CB-(i+1)I)L^{-1}\Lambda^{j}_{i+1}+L^{-1}\Lambda^{j}_{i+1}$
$=(CB-iI)L^{-1}\Lambda^{j}_{i+1}=\Lambda^{j}_{i}.\qed$
Lemma 0.
For $j\geq 1$, $\theta^{j}(I)=BL^{-1}\Lambda^{j}_{1}C$.
Proof.
For $j=1$, $\theta(I)=T=BL^{-1}C$, so the result holds. Suppose the result holds for $j\geq 1$. Then, $\theta^{j+1}(I)=(T+\partial_{x})(\theta^{j}(I))=BL^{-1}CBL^{-1}\Lambda_{1}^{j}C-BL^{-2}\Lambda_{1}^{j}C+BL^{-1}(\Lambda^{j}_{1})^{\prime}C$. By Lemma 3,
$\theta^{j+1}(I)=BL^{-1}\left((CB-I)L^{-1}\Lambda_{1}^{j}+(\Lambda^{j}_{1})^{\prime}\right)C=BL^{-1}\Lambda^{j+1}_{1}C.\qed$
Now by Lemmas 2 and 4, for $s\geq 0$, one can write
$\theta^{s}a=\sum_{j=0}^{s}\theta^{j}(I)\binom{s}{j}a^{(s-j)}=a^{(s)}+BL^{-1}\left[\sum_{j=1}^{s}\Lambda^{j}_{1}C\binom{s}{j}a^{(s-j)}\right].$
For $i\geq 1$ and $s\geq 0$, let $u_{i}^{s}=C\binom{s}{i}a^{(s-i)}$ and $\lambda_{i}^{s}=\sum_{j=i}^{s}\Lambda_{i}^{j}u_{j}^{s}$. Note that $u_{i}^{s}=0$ and $\lambda_{i}^{s}=0$ when $i>s$. Also $u_{i}^{s}$ is a vector of polynomials. So we have $\theta^{s}a=a^{(s)}+BL^{-1}\lambda_{1}^{s}$. Hence the matrix $K$ in (11) can be written as
$K=\begin{bmatrix}a^{(s_{1})}&\cdots&a^{(s_{r})}\end{bmatrix}+BL^{-1}\mathcal{L}_{1},$
with $\mathcal{L}_{i}=\begin{bmatrix}\lambda_{i}^{s_{1}}&\cdots&\lambda_{i}^{s_{r}}\end{bmatrix}$ for $i\geq 1$. In the above sum, the left term is a polynomial matrix. So by Proposition 3, $\varphi_{\ell}(K)$ divides $\varphi_{\ell}(L^{-1})\varphi_{\ell}(\mathcal{L}_{1})$. Therefore $\varphi_{\ell}(K)$ divides $\Delta\varphi_{\ell}(\mathcal{L}_{1})$. It only remains to prove that $\varphi_{\ell}(\mathcal{L}_{1})$ divides $\Delta^{s_{r}-1}$. This is done in the following lemma, which completes the proof of Proposition 1.
Lemma 0.
For $1\leq i\leq s_{r}$ and $\ell\geq 0$, $\varphi_{\ell}(\mathcal{L}_{i})$ divides $\Delta^{s_{r}-i}$.
Proof.
$\mathcal{L}_{s_{r}}$ is a polynomial matrix, so $\varphi_{\ell}(\mathcal{L}_{s_{r}})=1$ for all $\ell\geq 0$. Suppose the result holds for $1<i+1\leq s_{r}$. For all $s\geq 0$, we have
$\lambda_{i}^{s}=\sum_{j=i}^{s}\Lambda_{i}^{j}u_{j}^{s}=u_{i}^{s}+\sum_{j=i+1}^{s}\Lambda_{i}^{j}u_{j}^{s}=u_{i}^{s}+(CB-iI)L^{-1}\lambda_{i+1}^{s}.$
Thus, one can write
$\mathcal{L}_{i}=\begin{bmatrix}u_{i}^{s_{1}}&\cdots&u_{i}^{s_{r}}\end{bmatrix}+(CB-iI)L^{-1}\mathcal{L}_{i+1}.$
Since the left term is a polynomial matrix, by Proposition 3, $\varphi_{\ell}(\mathcal{L}_{i})$ divides $\varphi_{\ell}(L^{-1})\varphi_{\ell}(\mathcal{L}_{i+1})$. Hence by induction it divides $\Delta^{s_{r}-i}$. ∎
5.2. Proof of the degree bound
We are now ready to establish the degree bound stated in Theorem 1. The beginning of the proof does not differ from the proof of Theorem 1, except that we take $\Delta$ as a denominator of $T=XM^{-1}Y$.
Proof of Theorem 1.
We have $T=XM^{-1}Y$ so one can write $T$ with denominator $\Delta=\det M$. And because $T$ is strictly proper, $\Delta T$ is a polynomial matrix of degree $<\delta=\deg\det M$. Thus similarly as in the proof of Theorem 1 in Section 3, one can write $\theta^{i}a$ for all $i\geq 0$ as $b_{i}/\Delta^{i}$ with $b_{i}\in k[x]^{n}$ of degree at most $d_{a}+i(\delta-1)$.
Let us again study the solution $\nu\in k(x)^{\rho}$ of the linear system $K\nu=-\theta^{\rho}a$ with $K=\begin{bmatrix}a&\theta a&\cdots&\theta^{\rho-1}a\end{bmatrix}\in k(x)^{n\times\rho}$. We perform a valid deletion of $(n-\rho)$ rows in this system to make it square and obtain $\tilde{K}\nu=c$ where $\tilde{K}$ and $c$ are obtained after deleting the corresponding rows in $K$ and $-\theta^{\rho}a$.
Therefore by Cramer’s rule, we can still write $\nu_{i}=\det K_{i}/\det\tilde{K}$ for $0\leq i<\rho$ where $K_{i}$ is obtained from $\tilde{K}$ after replacing column $i+1$ by $c$. By a direct expansion, we get $\det\tilde{K}=p/\Delta^{\rho(\rho-1)/2}$ with $p\in k[x]$ of degree at most $\rho d_{a}+1/2\cdot\rho(\rho-1)(\delta-1)$. However, $\det\tilde{K}$ is a minor of $K$, thus by Proposition 1, it can be written with denominator $\Delta^{\rho-1}$. Thus we have
$\det\tilde{K}=\frac{p}{\Delta^{\rho(\rho-1)/2}}=\frac{q}{\Delta^{\rho-1}},$
with $q=p/\Delta^{(\rho-1)(\rho-2)/2}$ a polynomial of degree at most
$\deg q\leq\rho d_{a}+(\rho-1)\delta-1/2\cdot\rho(\rho-1).$
The same situation occurs for $\det K_{i}$. Indeed it can be written $p_{i}/\Delta^{\rho(\rho+1)/2-i}$ with $p_{i}\in k[x]$ of degree at most $\rho d_{a}+(1/2\cdot\rho(\rho+1)-i)(\delta-1)$. But up to a sign, $\det K_{i}$ can be seen as a minor of the matrix $K^{*}=\begin{bmatrix}a&\cdots&\theta^{\rho-1}a&\theta^{\rho}a\end{bmatrix}$. Hence by Proposition 1, it can be written with denominator $\Delta^{\rho}$. So we have
$\det K_{i}=\frac{p_{i}}{\Delta^{\rho(\rho+1)/2-i}}=\frac{q_{i}}{\Delta^{\rho}},$
with $q_{i}=p_{i}/\Delta^{\rho(\rho-1)/2-i}$ a polynomial of degree at most
$\deg q_{i}\leq\rho d_{a}+\rho\delta-(1/2\cdot\rho(\rho+1)-i).$
Finally for $0\leq i<\rho$, $\nu_{i}=q_{i}/(\Delta q)$. And thus $\eta_{i}=\Delta q\cdot\nu_{i}=q_{i}$ for $0\leq i<\rho$ and $\eta_{\rho}=\Delta q$, whence the result. ∎
6. Perspectives
Theorem 1 allows to establish precise degree bounds in several concrete problems under some conditions (genericity, point at infinity not irregular). For future work, we aim at proving a similar result without the strict properness assumption. This would yield bounds that hold unconditionally in our four instances. Also, there are other problems that are instances of Problem 1 for which we could also derive bounds. This is for example the case for the composition of algebraic functions and $D$-finite functions (kauers2017bounds_algsubs, ). Finally, the long-term goal is to exploit the unified viewpoint we enlighted for the design of new efficient algorithms. We notably intend to have algorithms solving Problem 1 whose runtime is sensitive to the size of the realisation of $T$ we have in input.
Acknowledgements.
The author thanks Alin Bostan, Bruno Salvy and Gilles Villard for helpful discussions and comments. This work has been supported by the French–Austrian project EAGLES (ANR-22-CE91-0007 & FWF I6130-N) and Agence nationale de la recherche (ANR) project NuSCAP (ANR-20-CE48-0014).